OpenStudy (anonymous):

two numbers total 87 and have a difference of 11. What are they?

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@Pamela16

OpenStudy (anonymous):

Any answer choices?

OpenStudy (anonymous):

nevermind it's 49 and 38!

OpenStudy (danjs):

You could set up 2 equations, with two variables as your unknown numbers you are trying to find.

OpenStudy (anonymous):

Oh! Yay, you found the answer!! XD

OpenStudy (danjs):

x = first number y = second number The sum is 87 x + y = 87 The difference is 11 x - y = 11 The two equations to solve are x + y = 87 x - y = 11